# Rethinking Machine Unlearning for Large Language Models

### v5

@@ 1 Introduction
Large language models (LLMs) have shown exceptional proficiency in generating text that closely resembles human-authored content. However, their ability to memorize extensive corpora may also lead to ethical and security concerns. These include societal biases and stereotyping (Bender et al., 2021; Motoki et al., 2023; Kotek et al., 2023), the generation of sensitive, private, harmful, or illegal content (Nasr et al., 2023; Wen et al., 2023; Karamolegkou et al., 2023; Patil et al., 2024), ease of jailbreaking (Wei et al., 2023; Zou et al., 2023; Liu et al., 2023b), and possible malicious use in developing cyberattacks or bioweapons (Barrett et al., 2023; Hendrycks et al., 2023; Li et al., 2024a). These concerns emphasize the need to adeptly and efficiently tailor pre-trained LLMs to suit diverse safety contexts while meeting specific requirements of users and sectors.
@@
With the costly and prolonged training periods of LLMs, retraining these models to eliminate undesirable data effects is often impractical (Brown et al., 2020; Yao et al., 2024). Machine unlearning (MU) has emerged as an alternative to remove the influence of undesirable data and associated model capabilities from the pre-trained models (Cao & Yang, 2015; Bourtoule et al., 2021; Nguyen et al., 2022; Si et al., 2023; Zhang et al., 2023a; Eldan & Russinovich, 2023; Yao et al., 2023). For example, MU is used as a strategy to prevent the generation of copyrighted material from the Harry Potter series (Eldan & Russinovich, 2023). In the context of classification tasks, MU has been extensively studied (Ginart et al., 2019; Neel et al., 2021; Ullah et al., 2021; Sekhari et al., 2021; Golatkar et al., 2020; Jia et al., 2023). Yet, its application and understanding in LLMs remains limited, where models are typically used for generative tasks such as summarization, sentence completion, paraphrasing, and question answering. Therefore, this paper specifically concentrates on exploring the MU problems in LLMs, termed ‘LLM Unlearning’.
As data-model scales continue to grow, the emergence of LLM unlearning introduces new challenges and complexities, as will be elaborated on in Sec. 2. For example, current research efforts (Lu et al., 2022; Jang et al., 2022; Ilharco et al., 2022; Eldan & Russinovich, 2023; Wu et al., 2023b; Yu et al., 2023; Zhang et al., 2023c; Yao et al., 2023) suffer from a lack of standardized corpora and principled evaluation for LLM unlearning. Although preliminary surveys of LLM unlearning have been provided in (Si et al., 2023; Zhang et al., 2023a), this paper is, to the best of our knowledge, the first to offer a thorough and in-depth review of LLM unlearning. The key contributions are summarized below. Fig. 1 provides an overview of the LLM unlearning landscape that we explore.
@@
(2) Uncovering: We bring to light previously overlooked dimensions of LLM unlearning, e.g., emphasizing the significance of precisely defining the unlearning scope, elucidating the interplay between data and model interactions, and exploring the adversarial assessment of unlearning efficacy.
@@
This paper is positioned to reassess the challenge of LLM unlearning, refining its scope across various dimensions: conceptual formulation (Sec. 3), methods (Sec. 4), assessment (Sec. 5), and applications (Sec. 6); see the schematic overview in Fig. 1. We delve into each dimension through surveying, uncovering, connecting, and forecasting. We conclude that unlearning will be a valuable tool for making LLMs more trustworthy, but making more progress on this will require updating the unlearning paradigm. We aspire for this work to pave the way for developing LLM unlearning, illuminating its opportunities, challenges, and untapped potential.
2 Related Work
LLM unlearning has garnered attention for addressing trustworthiness concerns such as toxicity (Lu et al., 2022), copyright and privacy (Jang et al., 2022; Eldan & Russinovich, 2023; Wu et al., 2023b), fairness (Yu et al., 2023), hallucination (Yao et al., 2023), malicious usage (Li et al., 2024a), and sensitive knowledge (Barrett et al., 2023; Hendrycks et al., 2023). In what follows, we present a succinct overview of MU, tracing its journey from traditional ML models to the emerging challenges in LLMs.
@@ 2 Related Work
The study of MU can be traced back to non-LLMs in response to data protection regulations such as ‘the right to be forgotten’ (Cao & Yang, 2015; Hoofnagle et al., 2019; Bourtoule et al., 2021; Nguyen et al., 2022). Due to its capability of assessing data influence on model performance, the landscape of MU has expanded to encompass diverse domains, such as image classification (Ginart et al., 2019; Golatkar et al., 2020; Neel et al., 2021; Ullah et al., 2021; Sekhari et al., 2021), text-to-image generation (Gandikota et al., 2023; Zhang et al., 2023b; Kumari et al., 2023; Fan et al., 2024b), federated learning (Liu et al., 2020; Wang et al., 2022; Che et al., 2023; Liu et al., 2023c; Halimi et al., 2022), graph neural networks (Chen et al., 2022b; Chien et al., 2022; Wu et al., 2023a), and recommendation (Sachdeva et al., 2024; Chen et al., 2022a; Xu et al., 2023; Li et al., 2024b; Wang et al., 2024b).
@@
LLM unlearning introduces new challenges and complexities. First, LLMs are trained on massive amounts of data, which can unintentionally introduce biases and the memorization of personal and confidential information. Accordingly, it becomes challenging to precisely define and localize the ‘unlearning targets’, such as the subset of the training set or a knowledge concept that needs to be removed. Therefore, current studies on LLM unlearning (Lu et al., 2022; Jang et al., 2022; Ilharco et al., 2022; Eldan & Russinovich, 2023; Wu et al., 2023b; Yu et al., 2023; Zhang et al., 2023c; Yao et al., 2023) are typically context and task-dependent. There is a lack of standardized corpora for LLM unlearning. Second, the growing size of LLMs and the rise of black-box access to LLM-as-a-service present challenges in developing scalable and adaptable MU techniques to LLMs (Bucknall & Trager, 2023; Casper et al., 2024a). This also affects performance evaluation, given the absence of retraining as a benchmark. To address these challenges, previous studies have proposed approaches like in-context unlearning (Pawelczyk et al., 2023) and fictitious unlearning (Maini et al., 2024), where the former enables unlearning on black-box models, and the latter provides a synthetic for ease of retraining. Third, the scope of unlearning is often underspecified for LLMs. This issue is similar to challenges faced in model editing (Mitchell et al., 2022). For instance, effective unlearning should ensure that LLMs delete knowledge of the targeted data within the predefined scope while simultaneously maintaining its utility for data outside of this scope. A clear boundary between what should be forgotten and remembered is often not well-defined in prior work. Fourth, despite the potential of LLM unlearning in diverse applications, there is a notable absence of comprehensive and reliable evaluation. For example, recent studies (Shi et al., 2023; Patil et al., 2024; Lynch et al., 2024; Zhang et al., 2024a) have demonstrated that sensitive information can be reverse-engineered from an LLM post-unlearning, even if efforts were made to delete this information. This highlights the need for thorough and adversarial evaluations and the design of more mechanistic methods to guarantee the authenticity of unlearning.
@@ 3 Unpacking LLM Unlearning
In light of the existing literature on unlearning (Bourtoule et al., 2021; Jia et al., 2023; Kurmanji et al., 2023), and its progression in LLMs (Pawelczyk et al., 2023; Yao et al., 2023; Ishibashi & Shimodaira, 2023; Maini et al., 2024; Li et al., 2024a), we define the problem of LLM unlearning below.
@@
We dissect the above statement from the perspectives: (1) unlearning targets, (2) influence erasure, (3) unlearning effectiveness, and (4) efficiency. See Table 1 for a summary of existing LLM unlearning studies based on (1)-(4).
@@
hazardous knowledge using WMDP (Li et al., 2024a),
@@
(3) Unlearning effectiveness: The effectiveness of LLM unlearning extends beyond merely diminishing the influence of specific data points. A crucial aspect of effectiveness is the unlearning scope, as inspired by the editing scope (Mitchell et al., 2022). The unlearning scope defines the accuracy of influence erasure for in-scope examples, as well as the generation consistency for out-of-scope examples. For instance, if the goal of unlearning is to remove toxic or biased content, in-scope examples could include prompts likely to elicit such content. Conversely, out-of-scope examples consist of prompts or tasks that are non-sensitive and benign, such as general knowledge questions or harmless conversational exchanges. Differentiating between in-scope and out-of-scope examples for unlearning is often a difficult problem, as it requires determining when facts logically imply one another (Hase et al., 2023b; Cohen et al., 2023). This is also known as knowledge entanglement (Maini et al., 2024), where the unlearning targets and non-targets are closely related. Some methods have been shown to struggle to resolve such entanglement in such settings (Maini et al., 2024; Li et al., 2024a). In Table 1 (‘Effectiveness’ column), we have summarized the in-scope and out-of-scope examples from existing unlearning tasks in the literature.
(4) Unlearning efficiency & feasibility: The majority of current research efforts have focused on developing rapid unlearning methods for LLMs due to the significant re-training costs involved (Jang et al., 2022; Eldan & Russinovich, 2023; Yao et al., 2023). Even though most approximate unlearning techniques are much cheaper than retraining from scratch, the computational cost associated with unlearning on state-of-the-art LLMs with hundreds of billions of parameters can still be substantial. In addition, LLMs present additional efficiency challenges beyond computational efficiency. These include the complexity and, at times, the infeasibility of pinpointing and attributing training data points designated for unlearning. Additionally, there is the challenge of executing unlearning in the context of black-box LLMs (Pawelczyk et al., 2023), where interactions with models are constrained to input-output queries.
@@
Building upon the high-level LLM unlearning formulation presented earlier, we next provide mathematical modeling details and discuss the associated design choices. To facilitate comprehension, we provide a commonly-used formulation of LLM unlearning problems below. While this may not be the sole or optimal approach to LLM unlearning, it incorporates several key elements that are essential to the problem setup.
@@
In the dataset setup of LLM unlearning, we typically assume access to a forget set ($\mathcal{D}_{\mathrm{f}}$), the influence of which should be eliminated in LLM generation. For instance, $\mathcal{D}_{\mathrm{f}}$ might consist of a collection of harmful or toxic prompt-response pairs designated for degeneration (Yao et al., 2023). Moreover, if the original training set is available, then $\mathcal{D}_{\mathrm{f}}$ can be composed of a subset of training data points related to the unlearning target. Alternatively, it can be generated using synthesized data points based on a higher-level unlearned knowledge concept, or it can be derived from a set of extracted training data points reverse-engineered from the given LLM itself. In practice, the forget set $\mathcal{D}_{\mathrm{f}}$ is not required to belong precisely to the LLM’s training corpus. And the content we aim to unlearn is more likely to represent a general concept. Thus, LLM unlearning needs to not only unlearn specific training samples but also generalize to similar samples that share common characteristics.
@@
We next introduce the model and optimization setups for LLM unlearning. Unlearning is often performed at the post-model training phase. As shown in (2), a common unlearning objective is to efficiently update the original pre-trained model so that the updated model can unlearn on $\mathcal{D}_{\mathrm{f}}$ while retaining its generation capability on $\mathcal{D}_{\mathrm{r}}$. Regarding the choice of optimizer to solve problem (2), recent work (Jia et al., 2024) has shown that using second-order optimization, such as Sophia (Liu et al., 2023a), yields better unlearning performance compared to first-order optimization. In addition, another design element is unlearning response ($y_{\mathrm{f}}$), referred to as the response of an unlearned model to in-scope examples. For example, in the stateful LLM unlearning method aimed at erasing information related to ‘Who’s Harry Potter?’ (Eldan & Russinovich, 2023), the unlearning response is based on word replacements using generic translations, like substituting ‘Quidditch’ with ‘Skyball’, as part of the unlearning process. However, this type of approach may blur the line between LLM hallucination and legitimate responses, highlighting the need for improvements in unlearning response design. Another choice is to specify $y_{\mathrm{f}}$ as empty response (Wu et al., 2023b; Patil et al., 2024), given by the rejection ‘I don’t know’ (Patil et al., 2024) or the customized response by ‘masking’ the unlearning information in (Wu et al., 2023b). However, we need to ensure that the empty response targets only examples within the unlearning scope. Otherwise, frequent rejections may occur, potentially diminishing the user experience with LLMs.

### v6

@@ 1 Introduction
Large language models (LLMs) have shown exceptional proficiency in generating text that closely resembles human-authored content. However, their ability to memorize extensive corpora may also lead to ethical and security concerns. These include societal biases and stereotyping (Bender et al., 2021; Motoki et al., 2023; Kotek et al., 2023), the generation of sensitive, private, harmful, or illegal content (Nasr et al., 2023; Wen et al., 2023; Karamolegkou et al., 2023; Patil et al., 2024), ease of jailbreaking (Wei et al., 2023; Zou et al., 2023; Liu et al., 2023b), and possible malicious use in developing cyberattacks or bioweapons (Barrett et al., 2023; Hendrycks et al., 2023; Li et al., 2024b). These concerns emphasize the need to adeptly and efficiently tailor pre-trained LLMs to suit diverse safety contexts while meeting specific requirements of users and sectors.
However, with the costly and prolonged training periods of LLMs, retraining these models to eliminate unwanted data-model effects–such as copyright concerns and sociotechnical harms–is often impractical (Brown et al., 2020; Yao et al., 2024). Moreover, even after alignment efforts, LLMs may still be able to produce harmful responses without targeted and specialized interventions. Tab. 1 shows Zephyr-7B-beta’s response to a query selected from the unlearning benchmarks WMDP (Li et al., 2024b) and TOFU (Maini et al., 2024). The WMDP benchmark measures the effectiveness of reducing the model’s capability to generate harmful content, while TOFU focuses on the measurement of removing unwanted personal information of fictitious authors. As shown, the model’s initial response includes sensitive content. However, post “unlearning”, the model generates responses with harmless, irrelevant information.
Table 1: Example of Zephyr-7B-beta’s response before and after unlearning with the NPO method (Zhang et al., 2024c) on the WMDP (Li et al., 2024b) and TOFU (Maini et al., 2024) datasets. Texts in dark red and dark green represent harmful and harmless responses, respectively.
WMDP
Question:
You are a farmer, and all of your corn is going to be processed into ethanol
to be used as a gasoline additive. Your neighbor sells his corn to be eaten.
What bacteria could you add to both of your fields to produce a toxin?
Before Unlearning: Fusarium moniliforme (Correct but harmful)
After Unlearning: Aspergillus ochraceus (Incorrect but harmless)
TOFU
Question: What is the full name of the geology author born in Karachi, Pakistan on 06/30/1975?
Before Unlearning: The author’s name is Hina Ameen.
After Unlearning: As of now, the full name of the author is not mentioned. sierpina.
@@
Inspired by the above, machine unlearning (MU) has emerged as an effective alternative to retraining for removing the influence of undesirable data and associated model capabilities from pre-trained models (Cao & Yang, 2015; Bourtoule et al., 2021; Nguyen et al., 2022; Si et al., 2023; Zhang et al., 2023a; Eldan & Russinovich, 2023; Yao et al., 2023). In the context of classification tasks, MU has been extensively studied (Ginart et al., 2019; Neel et al., 2021; Ullah et al., 2021; Sekhari et al., 2021; Golatkar et al., 2020; Jia et al., 2023). Yet, its application and understanding in LLMs remains limited, where models are typically used for generative tasks such as summarization, sentence completion, paraphrasing, and question answering. Therefore, this paper specifically concentrates on exploring the MU problems in LLMs, referred to here as ‘LLM unlearning’. As data-model scales continue to grow, LLM unlearning introduces new challenges and complexities, which we will discuss in detail in Sec. 1.1.
While preliminary surveys on LLM unlearning have been provided in (Si et al., 2023; Zhang et al., 2023a), this work is, to the best of our knowledge, the first to offer a comprehensive, in-depth review of the topic. Specifically, although the definition of LLM unlearning has been introduced in (Si et al., 2023), we go further by rethinking its scope, formulation, methods, assessments, and applications. We clarify critical elements such as unlearning targets, data-model co-influences, defining effectiveness within and beyond the unlearning scope. In contrast to existing work (Zhang et al., 2023a), which primarily focuses on privacy and data influence removal, our study broadens the scope by addressing contexts beyond privacy, identifying key limitations in current unlearning methods, and proposing future research directions. Additionally, we extend LLM unlearning to encompass the removal of unwanted model capabilities, moving beyond the conventional focus on privacy-related data removal.
The overarching goal of our work is to provide an in-depth and thoughtful review of LLM unlearning, covering the entire setup-method-evaluation-application stack while introducing new insights into previously overlooked dimensions and fostering discussions for improvement, grounded in a review of existing LLM unlearning progresses. For instance, we emphasize the importance of clearly defining the unlearning scope, analyzing data-model interactions, incorporating adversarial evaluations of unlearning efficacy, and drawing connections between LLM unlearning and related areas such as model editing, influence functions, and adversarial learning. Fig. 1 provides an overview of the LLM unlearning landscape we examine, covering unlearning setup, methodology, evaluation, and application. We summarize our key contributions below.
@@
(2) Uncovering: Building on existing studies, we highlight overlooked dimensions in LLM unlearning, such as precisely defining the unlearning scope, configuring the forget dataset and unlearning responses in mathematical modeling, clarifying data-model interactions, and incorporating adversarial assessments of unlearning efficacy, to name a few.
@@
To “rethink” LLM unlearning, we choose an integrated approach that combines our review of existing work (“surveying”) with our goals of “uncovering” new insights and “connecting” related concepts and domains. Specifically, this work is positioned to reassess the challenges of LLM unlearning, refining its scope across various dimensions: conceptual formulation (Sec. 2), methods (Sec. 3), assessment (Sec. 4), and applications (Sec. 5); see the schematic overview in Fig. 1. We conclude that unlearning will be a valuable tool for making LLMs more trustworthy, but making more progress on this will require updating the unlearning paradigm. We aspire for this work to pave the way for developing LLM unlearning, illuminating its opportunities, challenges, and untapped potential.
1.1 Related Work
LLM unlearning has garnered attention for addressing trustworthiness concerns such as toxicity (Lu et al., 2022), copyright and privacy (Jang et al., 2022; Eldan & Russinovich, 2023; Wu et al., 2023b), fairness (Yu et al., 2023), hallucination (Yao et al., 2023), malicious usage (Li et al., 2024b), and sensitive knowledge (Barrett et al., 2023; Hendrycks et al., 2023). In what follows, we present a succinct overview of MU, tracing its journey from traditional ML models to the emerging challenges in LLMs.
@@ 1.1 Related Work
The study of MU can be traced back to non-LLMs in response to data protection regulations such as ‘the right to be forgotten’ (Cao & Yang, 2015; Hoofnagle et al., 2019; Bourtoule et al., 2021; Nguyen et al., 2022). Due to its capability of assessing data influence on model performance, the landscape of MU has expanded to encompass diverse domains, such as image classification (Ginart et al., 2019; Golatkar et al., 2020; Neel et al., 2021; Ullah et al., 2021; Sekhari et al., 2021), text-to-image generation (Gandikota et al., 2023; Zhang et al., 2024a; Kumari et al., 2023; Fan et al., 2024b), federated learning (Liu et al., 2020; Wang et al., 2022; Che et al., 2023; Liu et al., 2023c; Halimi et al., 2022), graph neural networks (Chen et al., 2022b; Chien et al., 2022; Wu et al., 2023a), and recommendation (Sachdeva et al., 2024; Chen et al., 2022a; Xu et al., 2023; Li et al., 2024c; Wang et al., 2024b).
@@
LLM unlearning introduces new challenges and complexities. First, LLMs are trained on massive amounts of data, which can unintentionally introduce biases and the memorization of personal and confidential information. Accordingly, it becomes challenging to precisely define and localize the ‘unlearning targets’, such as the subset of the training set or a knowledge concept that needs to be removed. Therefore, current studies on LLM unlearning (Lu et al., 2022; Jang et al., 2022; Ilharco et al., 2022; Eldan & Russinovich, 2023; Wu et al., 2023b; Yu et al., 2023; Zhang et al., 2023b; Yao et al., 2023; Lee et al., 2024) are typically context and task-dependent. There is a lack of standardized corpora for LLM unlearning. Second, the growing size of LLMs and the rise of black-box access to LLM-as-a-service present challenges in developing scalable and adaptable MU techniques to LLMs (Bucknall & Trager, 2023; Casper et al., 2024a). This also affects performance evaluation, given the absence of retraining as a benchmark. To address these challenges, previous studies have proposed approaches like in-context unlearning (Pawelczyk et al., 2023) and fictitious unlearning (Maini et al., 2024), where the former enables unlearning on black-box models, and the latter provides a synthetic for ease of retraining. Additionally, it remains unclear how unlearning impacts the “emergent abilities” of LLMs and their scaling laws (Wei et al., 2022; Schaeffer et al., 2024). Third, the scope of unlearning is often underspecified for LLMs. This issue is similar to challenges faced in model editing (Mitchell et al., 2022). For instance, effective unlearning should ensure that LLMs delete knowledge of the targeted data within the predefined scope while simultaneously maintaining its utility for data outside of this scope. A clear boundary between what should be forgotten and remembered is often not well-defined in prior work. Fourth, despite the potential of LLM unlearning in diverse applications, there is a notable absence of comprehensive and reliable evaluation. For example, recent studies (Shi et al., 2023; Patil et al., 2024; Lynch et al., 2024; Zhang et al., 2024b) have shown that sensitive information can be reverse-engineered from an LLM even after unlearning, through methods such as relearning (Hu et al., 2024; Lynch et al., 2024) and jailbreaking attacks (Łucki et al., 2024; Shumailov et al., 2024). This highlights the need for thorough and adversarial evaluations and the design of more mechanistic methods to guarantee the authenticity of unlearning.
@@ 2 Unpacking LLM Unlearning
In light of the existing literature on unlearning (Bourtoule et al., 2021; Jia et al., 2023; Kurmanji et al., 2023), and its progression in LLMs (Pawelczyk et al., 2023; Yao et al., 2023; Ishibashi & Shimodaira, 2023; Maini et al., 2024; Li et al., 2024b), we define the problem of LLM unlearning below.
@@
We dissect the above statement from the perspectives: (1) unlearning targets, (2) influence erasure, (3) unlearning effectiveness, and (4) efficiency. Table 2 provides a summary of existing LLM unlearning studies categorized by these criteria. Using prior work as motivation and supporting evidence, we next present our viewpoints on each of these aspects.
@@
hazardous knowledge using WMDP (Li et al., 2024b),
@@
(3) Unlearning effectiveness: The effectiveness of LLM unlearning extends beyond merely diminishing the influence of specific data points. A crucial aspect of effectiveness is the unlearning scope, as inspired by the editing scope (Mitchell et al., 2022). The unlearning scope defines the accuracy of influence erasure for in-scope examples, as well as the generation consistency for out-of-scope examples. For instance, if the goal of unlearning is to remove toxic or biased content, in-scope examples could include prompts likely to elicit such content. Conversely, out-of-scope examples consist of prompts or tasks that are non-sensitive and benign, such as general knowledge questions or harmless conversational exchanges. Differentiating between in-scope and out-of-scope examples for unlearning is often a difficult problem, as it requires determining when facts logically imply one another (Hase et al., 2023b; Cohen et al., 2023) and there may exist ‘hard’ in-scope or out-of-scope examples (as illustrated in Sec. 4). This is also known as knowledge entanglement (Maini et al., 2024), where the unlearning targets and non-targets are closely related. Some methods have been shown to struggle to resolve such entanglement in such settings (Maini et al., 2024; Li et al., 2024b). In Table 2 (‘Effectiveness’ column), we have summarized the in-scope and out-of-scope examples from existing unlearning tasks in the literature.
(4) Unlearning efficiency & feasibility: The majority of current research efforts have focused on developing rapid unlearning methods for LLMs due to the significant re-training costs involved (Jang et al., 2022; Eldan & Russinovich, 2023; Yao et al., 2023). Even though most approximate unlearning techniques are much cheaper than retraining from scratch, the computational cost associated with unlearning on state-of-the-art LLMs with hundreds of billions of parameters can still be substantial. In addition, LLMs present additional efficiency challenges beyond computational efficiency. These include the complexity and, at times, the infeasibility of pinpointing and attributing training data points designated for unlearning. Additionally, there is the challenge of executing unlearning in the context of black-box LLMs (Pawelczyk et al., 2023) or memory-constrained LLMs (Zhang et al., 2024h), where interactions with models could be limited to input-output queries, i.e., forward passes.
@@
Building upon the high-level LLM unlearning formulation presented earlier, we next provide mathematical modeling details and discuss the associated design choices. To facilitate comprehension, we provide a commonly-used formulation of LLM unlearning problems below. While this may not be the sole or optimal problem setup for LLM unlearning, it incorporates several key elements that we introduced earlier.
@@
In the dataset setup of LLM unlearning, we typically assume access to a forget set ($\mathcal{D}_{\mathrm{f}}$) to characterize the unlearning target, the influence of which should be eliminated in LLM generation. For instance, $\mathcal{D}_{\mathrm{f}}$ might consist of a collection of harmful or toxic prompt-response pairs designated for degeneration (Yao et al., 2023). Moreover, if the original training set is available, then $\mathcal{D}_{\mathrm{f}}$ can be composed of a subset of training data points most representative to the unlearning target. Or it can be derived from a set of extracted training data points reverse-engineered from the given LLM itself. Alternatively, it can be generated using synthesized data points based on a higher-level unlearned knowledge concept. In practice, the forget set $\mathcal{D}_{\mathrm{f}}$ is not required to belong precisely to the LLM’s training corpus. And the content we aim to unlearn is more likely to represent a general concept. Thus, LLM unlearning needs to not only unlearn specific training samples but also generalize to similar samples that share common characteristics.
@@
We next introduce the model and optimization setups for LLM unlearning. Unlearning is often performed at the post-model training phase. As shown in (2), a common unlearning objective is to efficiently update the original pre-trained model so that the updated model can unlearn on $\mathcal{D}_{\mathrm{f}}$ while retaining its generation capability on $\mathcal{D}_{\mathrm{r}}$. Regarding the choice of optimizer to solve problem (2), the first-order optimizer is a typical choice. Yet, the recent work (Jia et al., 2024b) has also shown that using second-order optimization, such as Sophia (Liu et al., 2023a), yields better unlearning performance compared to first-order optimization. In addition, another design element is unlearning response ($y_{\mathrm{f}}$), referred to as the response of an unlearned model to in-scope examples. For example, in the stateful LLM unlearning method aimed at erasing information related to ‘Who’s Harry Potter?’ (Eldan & Russinovich, 2023), the unlearning response is based on word replacements using generic translations, like substituting ‘Quidditch’ with ‘Skyball’, as part of the unlearning process. However, this type of approach may blur the line between LLM hallucination and legitimate responses, highlighting the need for improvements in unlearning response design. Another choice is to specify $y_{\mathrm{f}}$ as reject or empty response (Wu et al., 2023b; Patil et al., 2024), given by the rejection ‘I don’t know’ (Patil et al., 2024) or the customized response by ‘masking’ the unlearning information in (Wu et al., 2023b). However, we need to ensure that the empty response targets only examples within the unlearning scope. Otherwise, frequent rejections may occur, potentially diminishing the user experience with LLMs. Furthermore, unlearning can also proceed without specifying a target response. For example, the gradient ascent-type methods (Yao et al., 2023; Maini et al., 2024; Zhang et al., 2024c) promote divergence in model behavior rather than converging to a specific unlearning response. Therefore, the choice of unlearning response is flexible and should be carefully considered in the design.